# hep-ph/***

### v3

@@ Abstract
Nilpotent deformations of an arbitrary background spacetime are considered. Those are of the form $g_{\mu\nu}+l_{\mu}l_{\nu}$ with $l^{2}=0$. The relationship between the Ricci tensor of the background metric and the Ricci tensor of the deformed metric is found exactly. It consists of two terms, one is essentially the Fierz-Pauli operator, and the other is new. When the background is flat, the Kerr-Schild family is recovered. Novel examples for more general backgrounds (even including some simple sources) are discussed.
@@ 1 Introduction
is not a solution. This is perhaps the main fact that prevents standard techniques (such as Fourier analysis) to be useful in this case.
@@
In this paper we are going to exploit a technique to generate new Ricci at spacetimes out of other ones. The idea is a generalization of the Kerr-Schild approach [7] and originated in [5] and [6].
@@
Next, let us compute an explicit and exact expressions of the Ricci tensor. Cartan’s connection can be easily computed as
@@
The proof is straightforward algebra. To be specific, let us give details on the simplest case. Assume that the perturbation is a Kerr-Schild metric
@@
where $l_{\mu}$ is a nilpotent vector, $l^{2}=0$, in which case (2) is trivially satisfied. Throughout this work we employ the standard optical variables, namely, the acceleration
@@
    We were able to generate new Ricci flat spacetimes out of a Ricci flat seed.
@@
    which is Ricci flat, when $H^{a}_{a}[u]=0$.
@@ 3 Sources
now the perturbation $h_{\mu\nu}$ induces a nonvanishing $R^{\text{\tiny{(1)}}}_{\alpha\beta}$ tensor, namely
@@
has the one-loop and two-loops counterterm zero up to $O(\kappa^{4})$.
@@ 5 Conclusions
with $l^{2}=0$. A general formula for the ensuing Ricci tensor has been obtained, namely.

### v5

@@ Abstract
Generalized Kerr-Schild nilpotent deformations of an arbitrary background spacetime $\overline{g}_{\mu\nu}(x)$ are considered. Those are of the form $\overline{g}_{\mu\nu}+l_{\mu}l_{\nu}$ with $l^{2}=0$. The relationship between the Ricci tensor of the background metric and the Ricci tensor of the deformed metric is found exactly. It consists of two terms, one is essentially the Fierz-Pauli operator, and the other is new. When the background is flat, the Kerr-Schild family is recovered. Novel examples for more general backgrounds (even including some simple sources) are discussed.
@@ 1 Introduction
is not a solution. This is perhaps the main fact that prevents standard techniques (such as Fourier analysis) to be useful in this case. When we talk about linear equations, we are assuming superposition in this sense.
@@
In this paper we are going to exploit a technique to generate new Ricci flat spacetimes out of some other Ricci flat spacetime, which we will dub the seed. The idea is a generalization of the approach in [7] and originated in [5] and [6].
In the original work of Kerr and Schild they examined Ricci flat metrics for which there is a gauge of the form
$g_{\mu\nu}=\eta_{\mu\nu}+l_{\mu}l_{\nu}$ (5)
(where $\eta_{\mu\nu}$ is the flat Minkowski metric) and they found all of those trough quite heavy calculations using a clever tretrad. In this paper by using an apparently coarser approach we shall find unexpected relationships with the flat space spin 2 Fierz-Pauli equation.
@@
Next, let us compute an explicit and exact expressions of the Ricci tensor. Cartan’s connection (that is the Christoffel connection corresponding to the full metric $g_{\mu\nu}$) can be easily computed as
@@
The proof is straightforward algebra.
To be specific, let us give details on the simplest case. Assume that the perturbation is a Kerr-Schild metric
@@
where $l_{\mu}$ is a nilpotent vector, $l^{2}=0$, in which case (2) is trivially satisfied. It has already been pointed out that in this case, and only in this case, the perturbative expansion has got only a finite number of terms.
Throughout this work we employ the standard optical variables, namely, the acceleration
@@
    We were able to generate new Ricci flat spacetimes out of a Ricci flat seed. All those spacetimes are closely related to the Kasner family. Consider the last case as an example. It lacks the isometry
    $z\rightarrow-z$ (38)
    In fact the spacetime metric can be written as
    $\text{d}s^{2}=2\left(\text{d}t+{1\over 2}t^{p_{2}}dz\right)^{2}-t^{2p_{1}}\text{d}x^{2}-t^{2p_{2}}\text{d}y^{2}-{1\over 2}t^{2p_{2}}dz^{2}$ (39)
@@
    which is Ricci flat, when $H^{a}_{a}[u]=0$. Those spacetimes again represent gravitational waves of a sort.
@@ 3 Sources
The resulting spacetime metric looks quite involved
$ds^{2}=2\,dt^{2}-{2\over 3}e^{2Ht}\left(dx^{2}+dy^{2}+dz^{2}\right)+{2\over\sqrt{3}}e^{Ht}\left(dtdx+dtdy+dtdz\right)+$
$+{2\over 3}e^{2Ht}\left(dxdy+dxdz+dydz\right)$ (51)
In spite of the appearances this cosmological metric shares rotational invariances with the seed spacetime.
The perturbation $h_{\mu\nu}$ now induces a nonvanishing $R^{\text{\tiny{(1)}}}_{\alpha\beta}$ tensor, namely
@@
has the one-loop and two-loops counterterm zero.
@@ 5 Conclusions
with $l^{2}=0$. Those are an extension of the original Kerr-Schild ansatz, which obviously covers the full space of metrics, just by starting with an appropiate seed metric.
A general formula for the ensuing Ricci tensor has been obtained, namely.